The perimeter of a rectangle is 64cm. The length is 4cm less than twice the width. Find the length and the width .

Answer by addingup(3677)

2L + 2W = 64
L = 2W - 4 substitute above for L:
2(2W - 4) + 2W = 64
4W - 8 + 2W = 64
6W = 72
W = 12
and L?
L = 2W - 4 = 2(12) - 4 = 24 - 4 = 20
---------------------------------------------
Check:
2L + 2W = 64
2(20) + 2(12) = 64
40 + 24 = 64 Correct